The most recent financial statements for Crosby, Inc., follow. Interest expense will remain constant; the tax rate and the dividend payout rate will also remain constant. Costs, other expenses, current assets, fixed assets, and accounts payable increase spontaneously with sales. Assume the firm is operating at full capacity and the debt- equity ratio is held constant. CROSBY, INC. 2017 Income Statement $763,000 619,000 29,000 Sales Costs Other expenses Earnings before interest and taxes $115,000 Interest paid 14,000 $101,000 25,250 Taxable income Taxes (25%) $ 75,750 Net income $37,140 Dividends Addition to retained 38,610 earnings CROSBY, INC. Balance Sheet as of December 31, 2017 Assets Liabilities and Owners' Equity Current liabilities Accounts payable Notes payable Current assets $ 61.400 25,240 34,660 Cash Accounts receivable 17,600 Total $79,000 Inventory 71,420 Long-term debt $111,000 Total 131,320 Owners' equity Common stock and paid-in $106.000 Fixed assets surplus Net plant and $220,000 Retained earnings 55,320 equipment Total $161,320 Total liabilities and owners $351,320 equity $351,320 Total assets Complete the pro forma income statements below. (Do not round intermediate calculations. Round your answers to the nearest whole dollar amount.)
LldDillies dnu OwIners cquIly ASseis Current liabilities Accounts payable Notes payable Current assets Cash Accounts receivable 25,240 34,660 61,400 17,600 $79,000 71,420 Inventory Total Total 131,320 $111,000 Long-term debt Owners' equity Common stock and paid-in $106.000 Fixed assets surplus Net plant and equipment $220,000 Retained earnings 55.320 $161,320 Total Total liabilities and owners' $351,320 Total assets 351,320 equity Complete the pro forma income statements below. (Do not round intermediate calculations. Round your answers to the nearest whole dollar amount.) Pro Forma Income Statement 15% Sales Growth 20% Sales Growth 45% Sales Growth Sales Costs Other expenses EBIT Interest paid Taxable income Taxes (25%) Net income Dividends Add to RE Calculate the EFN for 15, 20 and 45 percent growth rates. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answers to the nearest whole dollar amount.) 15% 20% 45% EFN
Answer #1
A B=A*1.15 (for Sales, cost & other expense) C=A*1.2 (for Sales, cost & other expense) D=A*1.45 (for Sales, cost & other expense)
2017 Income Statement 15% Sales growth 20% Sales growth 45% Sales growth
E Sales $763,000 $877,450 $915,600 $1,106,350
F Costs $619,000 $711,850 $742,800 $897,550
G Other expenses $29,000 $33,350 $34,800 $42,050
H=E-F-G EBIT $115,000 $132,250 $138,000 $166,750
I Interest paid $14,000 $14,000 $14,000 $14,000
J=H-I Taxable Income $101,000 $118,250 $124,000 $152,750
K=25%*J Taxes(25%) $25,250 $29,563 $31,000 $38,188
L=J-K Net Income $75,750 $88,687 $93,000 $114,563
M=L*0.490297 Dividend $37,140 $43,483 $45,598 $56,170
N=L-M Add to Retained Earnings $38,610 $45,204 $47,402 $58,393
Dividend Payout rate=37140/75750=      0.490297
Item Present Leve Percent of sales Projected Level(15% sales growth) Projected Level(20% sales growth) Projected Level(45% sales growth)
Sales $763,000 100.00% $877,450 $915,600 $1,106,350
Current asset $131,320 17.21% Current asset $151,018 17.21%*877450 $157,584 17.21%*915600 $190,414 17.21%*1106350
Fixed asset $220,000 28.83% Fixed asset $253,000 28.83%*877450 $264,000 28.83%*915600 $319,000 28.83%*1106350
Total Assets $351,320 46.04% Total Assets $404,018 $421,584 $509,414
Accounts Payable $61,400 8.05% Accounts Payable $70,610 8.05%*877450 $73,680 8.05%*915000 $89,030 8.05%*1106350
Notes Payable $17,600 2.31% Notes Payable $17,600 $17,600 $17,600
Totalcurrent liabilities $79,000 10.35% Total current liabilities $88,210 $91,280 $106,630
Long termDebt $111,000 14.55% Long termDebt $111,000 $111,000 $111,000
TotalLiabilities $190,000 24.90% TotalLiabilities $199,210 $202,280 $217,630
Owners Equity Owners Equity
Common Stock & paid in surplus $106,000 13.89% Common Stock & paid in surplus $106,000 $106,000 $106,000
Retained earnings $55,320 7.25% Retained earnings $100,524 (55320+45204) $102,722 (55320+47402) $113,713 (55320+58393)
Total Owners Equity $161,320 21.14% Total owners equity $206,524 $208,722 $219,713
Total Liabilities+Equity $351,320 46.04%
A Increase in Total Assets $52,698 (404018-351320) $70,264 (421584-351320) $158,094 (509414-351320)
B Increase in Accounts payable $9,210 (70610-61400) $12,280 (73680-61400) $27,630 (89030-61400)
C Add to retained earnings $45,204 $47,402 $58,393
D=A-B-C External Financial Need -$1,716 $10,582 $72,071
15% 20% 45%
EFN $0 $10,582 $72,071